# EvaGaussians: Event Stream Assisted Gaussian Splatting from Blurry Images

### v1

@@ 5 Conclusions
This paper introduces Event Stream Assisted Gaussian Splatting (EvaGaussians), a novel framework that seamlessly integrates the event streams captured by an event camera into the training of 3D-GS, effectively addressing the challenges of reconstructing high-quality 3D-GS from motion-blurred images. We conducted comprehensive experiments on two novel datasets, and the results demonstrate that our method outperforms previous state-of-the-art deblurring rendering techniques. Despite its promising performance, our method may still face challenges when reconstructing scenes with extremely intricate textures from severely blurred images. We will release our code and dataset for future research.
Appendix A Datasets
A.1 Synthetic Data Generation
We created 9 medium-scale and large-scale raw scenes using Blender[4], utilizing these scenes in our experiments. Additionally, we remade and redesigned the camera trajectories for the cozyroom, factory, pool, and tanabata scenes from DeblurNeRF[19].
A.1.1 Introduction of the Synthetic Scenes
In this work, we utilized Blender to create a comprehensive dataset featuring a variety of indoor and outdoor scenes. These scenes are depicted in Figure. 6, and Figure. 7. The scenes created are meticulously crafted to provide a realistic and challenging environment for testing and training.
Figure 6: Visualization of EvaGaussians-Blender Indoor Scenes. The sizes of the Cafe and Classroom scenes are approximately ${15\times 7\times 4}$ meters, while the Dormitory scene is approximately ${5\times 5\times 4}$ meters (with an additional outdoor garden, making the overall scene size ${20\times 20\times 6}$ meters).
@@ A.1.1 Introduction of the Synthetic Scenes
Figure 7: Visualization of EvaGaussians-Blender Outdoor Scenes. These scenes include rich details including skies, lakes, rivers, deserts, forests, stones, cities, roads, etc. All scenes cover an area of more than 1 square kilometer (relatively).
@@
Our dataset aims to provide a wide range of realistic and diverse environments to facilitate the development and testing of advanced algorithms in various fields. Each scene is carefully crafted to ensure high-quality and detailed representations, making this dataset a valuable resource for researchers and developers.
A.1.2 Camera Trajectory in Blender
(a) Desert
(b) Classroom
(c) Café
(d) Forests
(e) Lake
(f) Dormitory
(g) City Blocks
(h) London
Figure 8: Visualization of Camera Trajectory. The trajectories depicted were manually configured within Blender to ensure precise control over the camera paths. For the purpose of visualization and clarity, these trajectories have been normalized. This normalization process helps maintain consistent visibility and comparative analysis across different trajectory visualizations. Additionally, the normalization ensures that the visual representation accurately reflects the relative movements and orientations within the scene, providing a clear and comprehensive understanding of the camera dynamics throughout the depicted sequences.
A.1.3 Camera Settings
In our experiments conducted in Blender[4], we configured the virtual camera with a resolution of $400\times 600$ pixels for width and height, respectively, and set the scaling factor to 1.0. The virtual camera utilized a perspective model with a shutter speed of 1/180 seconds. Different focal lengths were employed for various scenes, and these parameters are documented in the corresponding data files.
A.1.4 Dataset Settings
We developed a dedicated script to generate motion blur and partition the dataset. Along a predefined virtual camera trajectory, we uniformly sampled 35 camera poses, adding a certain level of jitter to create the training set. We recorded the start and end times of the camera exposure, the positions, and the 20 intermediate frames during the exposure (obtained through linear interpolation between the start and end positions). We then uniformly sampled 100 camera poses along the same trajectory to form the test set. Using the event camera simulator from vid2e[7], we simulated the event data stream for the entire camera motion and synthesized the event bins from the events at the start and end of the exposure.
A.2 Real Data Capture
We use the Color DAVIS 346 event camera to record our real-time sequences and utilize the default camera settings provided in the DV software that comes with the camera.
@@ A.2.1 Camera Calibration
Figure 9: Visualization of Camera Calibration. Referencing [34], this figure illustrates the process and results of camera calibration using a checkerboard pattern. The left panel shows the checkerboard pattern captured from various positions and angles, with detected corner points utilized for calibration. The right panel presents the calibrated checkerboard pattern, demonstrating the corresponding points and lines between two cameras, which reflect the geometric relationship and accuracy achieved after calibration. Different colored lines indicate the correspondences between points during the calibration process.
@@
Figure 10: Visualization of Reprojection Errors and Epipolar Errors. The figure illustrates the 50 sets of reprojection errors and epipolar errors generated during the calibration process. The reprojection error $e_{r}$ represents the average discrepancy between the observed points and the projected points, calculated as shown in Equation equation 10. The epipolar error $e_{\text{epipolar}}$ represents the average distance between points in one camera and the epipolar lines calculated from the other camera for each pair of images, calculated as shown in Equation equation 11. As shown in the figure, the average reprojection error is approximately 0.5, and the average epipolar error is approximately 0.05, indicating a high level of accuracy in the calibration process.
@@ A.2.2 Camera Settings
We recorded the five real scenes using the calibration parameters obtained during the calibration process. By adjusting the indoor lighting and shooting angles, we ensured the richness of the recorded scene details. The event camera thresholds were set according to its official documentation, with a contrast positive threshold $C_{p}=0.143$ and a contrast negative threshold $C_{n}=0.225$. Additionally, the camera has a spatial resolution of 346 × 260 pixels, a temporal resolution of 1 $\mu$s, a typical latency of less than 1 ms, a maximum throughput of 12 MEps, and a dynamic range of approximately 120 dB (with 50% of the pixels responding to 80% contrast changes under 0.1-100k lux conditions). The contrast sensitivity is 14.3% (ON) and 22.5% (OFF) (with 50% of the pixels responding). These parameters ensure that the event camera can stably and efficiently record scene information under various lighting conditions and dynamic ranges.
Appendix B Additional Experiment Results
To further validate the performance of our proposed model, we provide supplementary experimental results in this section. The following subsections offer a detailed analysis of the deblurring performance on the object-level scenes from the EvaGaussians-Blender dataset and present visualization results of novel views and deblurring views for each scene (Section. B.1). Additionally, we quantitatively analyze the deblurring performance and visualize the deblurring view and novel synthesis results in the medium-scale scenes (Section. B.2), and present quantitative comparisons for deblurring view synthesis (DVS) in the large-scale scenes (Section. B.3). Furthermore, in Section. B.4, we provide detailed evaluation metrics for each scene in the EvaGaussians-Blender dataset and visualize novel view synthesis (NVS) renderings of custom scenes in Section. B.5. Finally, we present evaluation metrics for each scene in the EvaGaussians-DAVIS dataset in Section. B.6.
B.1 Results of DVS in Object-level Scenes of EvaGaussians-Blender
In this section, we analyze the performance of our method in object-level scenes. Table. 6 and Figure. 11 present the quantitative and qualitative results of various methods across six synthetic scene sequences under default settings, which are optimal for all methods, as described in Section. 4.3. From the qualitative results, it is evident that our method excels in reconstructing fine details and maintaining high color accuracy in both deblurring view synthesis and novel view synthesis, as previously observed in Section. 4.3. In terms of quantitative results, our method outperforms baseline methods in most scenes and performs comparably in the remaining scenes. This is shown in the novel view synthesis results in Table 3 and the deblurring view synthesis results in Table 6.
Table 6: Quantitative comparisons of blur view synthesis for object-level scenes. The results presented in the table are the averages derived from six synthetic scenes from NeRF, utilizing 100 views in accordance with the ${\textnormal{E}^{2}}$eRF[30] training process, and evaluate our method with sharp images of the same 100 views. We highlight the best-performing results in red and the second-best results in orange.
Novel View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
@@ B.1 Results of DVS in Object-level Scenes of EvaGaussians-Blender
B.2 Results of DVS in Medium-scale Scenes of EvaGaussians-Blender
In this section, we present the quantitative deblurring results for medium-scale scenes, along with the visualization results of all models on the restructured DeblurNeRF dataset [19]. These experiments were conducted using the default parameters of each model. As shown in Table. 7, our method outperforms all others across all scenes. Additionally, Figure. 12 provides a detailed analysis of the deblurring and novel view synthesis results in the cozyroom, pool, factory, tanabata scenes which are medium scenes in the EvaGaussian-Blender Dataset. It can be observed that our model performs slightly worse than the NeRF baseline method in sky and water pool scenes. As noted in [48], the generalization performance of our model for transparent materials and sky scenes is inferior to the NeRF baseline method. However, our model excels in texture detail compared to other methods. These results demonstrate the superior learning capacity and generalization performance of our model.
Table 7: Quantitative comparisons of deblurring view synthesis on the medium-scale scenes of our EvaGaussians-Blender dataset. The table reports the average performance across four synthetic scenes, each trained utilizing 35 views and evaluate our method with sharp images of the same views. The results unequivocally demonstrate that our method surpasses previous state-of-the-art(SOTA) approaches, achieving superior performance consistently across all metrics.
Blur View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
@@ B.2 Results of DVS in Medium-scale Scenes of EvaGaussians-Blender
B.3 Results of DVS in Large-scale Scenes of EvaGaussians-Blender
As a supplement to the synthetic experiments detailed in Section. 4.3, we also performed experiments on deblurring view synthesis using the same experimental setup. The results are presented in Table. 8. Consistent with the findings in Section. B.1, these results further confirm the effectiveness of our model. Additionally, the per-scene breakdown of the EvaGaussians-Blender Dataset results is shown in Table. 9, Table. 10, and Table. 11. Similar to the deblurring view synthesis results, our method consistently achieves the best performance across all scenes.
Table 8: Quantitative comparison of deblurring view synthesis on the large-scale scenes in our EvaGaussians-Blender dataset. The table reports the average performance across four synthetic scenes, each trained utilizing 35 views and evaluate our method with sharp images of the same views. The results unequivocally demonstrate that our method surpasses previous SOTA approaches, achieving superior performance consistently across all metrics.
Novel View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
@@ B.3 Results of DVS in Large-scale Scenes of EvaGaussians-Blender
B.4 Per-scene Breakdown for Novel View Synthesis in Medium and Large Scale Scenes
In this subsection, we present a detailed analysis of the novel view synthesis performance in medium and large scale scenes from the EvaGaussians-Blender dataset. Table. 9 summarizes the Peak Signal-to-Noise Ratio results and the NVS results demonstrate that our proposed method consistently outperforms other approaches across various scenes. Specifically, our model achieves superior PSNR values, indicating higher reconstruction quality and better visual fidelity. The detailed metrics for Structural Similarity Index and Learned Perceptual Image Patch Similarity further corroborate these findings in Table. 10 and Table. 11, showing that our model excels in maintaining structural integrity and perceptual quality in synthesized views.
In medium-scale scenes, our method exhibits robust performance, particularly in complex environments where maintaining detail and minimizing artifacts are challenging. This robustness is evident in scenes such as cozyroom and factory, where our method achieves significant improvements in both PSNR and SSIM, reflecting enhanced clarity and texture detail. For large-scale scenes, our model continues to demonstrate its efficacy, especially in handling large areas and diverse visual features. Scenes like desert and city blocks highlight the model’s capability to generalize across different scales, providing high-quality novel view synthesis with minimal perceptual discrepancies. These comprehensive results underscore the advanced learning and generalization capabilities of our model, setting a new benchmark for novel view synthesis in medium- and large-scale scenes.
@@ B.4 Per-scene Breakdown for Novel View Synthesis in Medium and Large Scale Scenes
Ours 0.1981 0.0948 0.2328 0.1889 0.1920 0.0756 0.1037 0.2156 0.2862 0.3092 0.2924 0.2691
B.5 Visualization of Qualitative Results for NVS in EvaGaussians-Blender Datasets
In this section, we present the qualitative results of novel view synthesis for custom scenes in the EvaGaussians-Blender dataset, such as desert and city blocks. These visualizations highlight our model’s ability to reconstruct fine details and maintain high color accuracy. By comparing our method with baseline approaches, we demonstrate the visual fidelity and robustness of our model across different scenes under the same experimental settings. The synthesized views shown in Figure. 13 and Figure.  14 further exemplify the superior performance of our method in terms of detail preservation and color consistency.
B.6 Results of Deblurring and Novel View Synthesis in EvaGaussians-DAVIS Datasets
In this section, we present a comprehensive analysis of the deblurring and novel view synthesis results on the EvaGaussians-DAVIS dataset. The evaluation metrics include BRISQUE, NIQE, PIQE, MetaIQA, and RankIQA, which quantitatively assess the quality of synthesized views across different scenes such as desk & chair, washroom, pokémon, pillow, and bag.
Our method consistently achieves the lowest BRISQUE, NIQE, and PIQE scores, indicating superior perceptual quality in the synthesized views. Specifically, as shown in Table. 12, our model achieves the best BRISQUE scores across all scenes, highlighting its ability to produce visually appealing and less distorted images.
For NIQE, as presented in Table. 13, our approach significantly outperforms the baselines, achieving the lowest average NIQE score. This demonstrates our method’s robustness in generating high-quality images with minimal perceptual artifacts, further confirmed by its performance across individual scenes.
In terms of PIQE, Table. 14 shows that our model again leads in performance, achieving the lowest PIQE scores, which correlates with higher perceived image quality. This underscores the effectiveness of our model in preserving image details and reducing noise.
Furthermore, our method excels in MetaIQA and RankIQA evaluations, as detailed in Tables. 15 and Table. 16, respectively. The highest MetaIQA scores and lowest RankIQA scores across most scenes affirm the overall superior visual quality and fidelity of our synthesized views compared to baseline models.
Overall, these results demonstrate the advanced learning capabilities and generalization performance of our model, particularly in handling complex scenes and maintaining high visual quality across diverse scenarios. The consistent performance across multiple evaluation metrics solidifies our approach as a leading solution for novel view synthesis in realistic environments.
@@ B.6 Results of Deblurring and Novel View Synthesis in EvaGaussians-DAVIS Datasets
Ours 4.1037 5.9088 5.2513 5.2363 4.8332 5.0667
Figure 11: Visualization of DVS and NVS of Object-level Scenes in the EvaGaussian-Blender Dataset. It illustrates the performance of various methods in DVS and NVS tasks across multiple synthetic scenes. The results demonstrate that our method consistently excels in reconstructing fine details and maintaining high color accuracy compared to other methods. Notably, although our method performs less well in scenes with sparse elements such as $ficus,mic$, it shows superior performance in terms of visual fidelity and artifact reduction, closely matching the ground truth (GT) images.
Figure 12: Visualization of DVS and NVS of Redesigned DeblurNeRF Scenes in the EvaGaussian-Blender Dataset. This figure showcases the performance of various methods in DVS and NVS tasks across multiple synthetic scenes. The results demonstrate that our method consistently excels in reconstructing fine details and maintaining high color accuracy compared to other methods. Notably, similar to object-level scenes, our method performs less effectively in highly reflective scenes such as pool. However, it still exhibits superior performance in terms of visual fidelity and artifact reduction, closely matching the ground truth (GT) images.
@@
Appendix C Broader Impact
Our proposed EvaGaussians utilizes event cameras to perform 3D-reconstruction from low-quality, blurred images. It has the potential to bring about both positive and negative societal impacts.

### v2

@@ 5 Conclusions
This paper introduces Event Stream Assisted Gaussian Splatting (EvaGaussians), a novel framework that seamlessly integrates the event streams captured by an event camera into the training of 3D-GS, effectively addressing the challenges of reconstructing high-quality 3D-GS from motion-blurred images. We contribute two novel datasets and conduct comprehensive experiments. The results demonstrate that our method outperforms previous state-of-the-art deblurring rendering techniques in terms of novel view synthesis quality, without sacrificing inference efficiency. Despite its promising performance, our method may still face challenges when reconstructing scenes with extremely intricate textures from severely blurred images. We will release our code and dataset for future research.
Supplementary Material
Appendix A Dataset Details
A.1 EvaGaussians-Blender Dataset
A.1.1 Dataset Overview
We use Blender [5] to craft nine indoor and outdoor 3D scenes, and further incorporate four 3D scenes from DeblurNeRF [24] and six 3D objects from the NeRF-Synthetic dataset [26] as our base scenes. We then design various camera trajectories to simulate motion-blurred images on these base scenes, and generate the corresponding event streams using V2E [9]. Visualization of our crafted scenes are shown in Figure. 5 and Figure. 8, and an overview of the design purposes of these crafted scenes is provided below:
Figure 5: Visualization of EvaGaussians-Blender Indoor Scenes. The sizes of the Café and Classroom scenes are approximately ${15\times 7\times 4}$ meters, while the Dormitory scene is approximately ${5\times 5\times 4}$ meters (with an additional outdoor garden, making the overall scene size ${20\times 20\times 6}$ meters).
Figure 6: Visualization of Reprojection Errors and Epipolar Errors. The figure illustrates the 50 sets of reprojection errors and epipolar errors generated during the calibration process. The reprojection error $e_{r}$ represents the average discrepancy between the observed points and the projected points, calculated as shown in Eq. 10. The epipolar error $e_{\text{epipolar}}$ represents the average distance between points in one camera and the epipolar lines calculated from the other camera for each pair of images, calculated as shown in Eq. 11. As shown in the figure, the average reprojection error is approximately 0.5, and the average epipolar error is approximately 0.05, indicating a high level of accuracy in the calibration process.
Figure 7: Visualization of pose accuracy in different level of motion blur.
Figure 8: Visualization of EvaGaussians-Blender Outdoor Scenes. These scenes include rich details ane diverse components like sky, lake, river, desert, forest, cities, roads. All scenes cover an area of more than 1 square kilometer.
@@ A.1.1 Dataset Overview
Figure 9: Visualization of Camera Trajectory. The trajectories depicted were manually configured within Blender [5] to ensure precise control over the camera paths. For the purpose of visualization, these trajectories have been normalized.
A.1.2 Camera Settings
To render the base scenes and simulate motion blur, we configure the virtual camera in Blender with a resolution of $400\times 600$, and set the scaling factor to $1.0$. The virtual camera utilized a perspective model with a shutter speed of $1/180$ seconds. Subsequently, we developed a dedicated script to generate camera trajectory and motion blur. An example is shown in Figure. 9. Along each predefined virtual camera trajectory, we uniformly sampled 35 camera poses, adding a certain level of jitter to create the training set. We recorded the start and end time of the camera exposure time, the positions, and 20 intermediate frames during the exposure time (obtained through linear interpolation between the start and end positions). We then uniformly sample 100 camera poses along the same trajectory to form the test set. Using the event camera simulator from V2E [9], we simulate the event stream for each camera trajectory and synthesize the event bins from the event stream at the start and end of the exposure time.
A.2 EvaGaussians-DAVIS Dataset
We use the color DAVIS346 event camera [1] to record our real-world event and RGB sequences and utilize the default camera settings provided in the DV software that comes with the camera. We name the five captured scenes as desk & chair, washroom, pokémon, pillow, and bag.
@@ A.2.1 Camera Calibration
Figure 10: Illustration of Camera Calibration. The left panel shows the checkerboard pattern captured from various positions and angles, with detected corner points utilized for calibration. The right panel presents the calibrated checkerboard pattern, demonstrating the corresponding points and lines between two cameras, which reflect the geometric relationship and accuracy achieved after calibration. Different colored lines indicate the correspondences between points during the calibration process.
@@ A.2.2 Camera Settings
We recorded the five real scenes using the calibration parameters obtained during the calibration process. By adjusting the indoor lighting and shooting angles, we ensured the richness of the recorded scene details. The adopted event camera has a spatial resolution of $346\times 260$, a temporal resolution of 1 $\mu$s, a typical latency of less than 1 ms, a maximum throughput of 12 MEps, and a dynamic range of approximately 120 $dB$ (with 50% of the pixels responding to 80% contrast changes under 0.1-100k lux conditions). The contrast sensitivity is 14.3% (ON) and 22.5% (OFF) (with 50% of the pixels responding). These parameters ensure that the event camera can stably and efficiently record scene information under various lighting conditions and dynamic ranges.
Table 5: Quantitative comparisons of DVS on object-level scenes. The results indicate that our method outperforms previous state-of-the-art approaches, consistently achieving better performance across all metrics.
Deblur View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
@@
Table 6: Quantitative comparisons of DVS on the medium-scale scenes. The results show that our method surpasses previous state-of-the-art approaches, achieving better performance consistently across all metrics.
Deblur View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
@@
Table 7: Quantitative comparison of DVS on large-scale scenes. The results demonstrate that our method consistently achieves better performance across all metrics.
Deblur View B-NeRF B-3DGS UFP-GS EDI-GS EFN-GS $\mathbf{E}^{2}$eRF BAD-NeRF BAD-GS EDNeRF Ours
@@
Ours 0.1927 0.0928 0.2311 0.1859 0.1873 0.0715 0.1023 0.2053 0.2835 0.2983 0.2857 0.2674
@@
Table 12: Robustness of pose optimization.
           ATE $\downarrow$            Mild Blur            Medium Blur            Strong Blur
           COLMAP poses            $0.0598\pm 0.0167$            $0.0783\pm 0.0245$            $0.0973\pm 0.0352$
           Optimized poses            $0.0461\pm 0.0129$            $0.0586\pm 0.0149$            $0.0677\pm 0.0208$
@@
      Ours       3.9369       5.6564       5.0867       5.1538       4.6422       4.8952
Figure 11: NVS results on the EvaGaussian-DAVIS dataset. The first column shows the blurry image used for training, and the following rows show the deblurring results of different methods. The results demonstrate that our method consistently excels in reconstructing fine details compared to other methods [30, 52, 3].
Figure 12: Visualization of DVS and NVS of Object-level Scenes in the EvaGaussian-Blender Dataset. The DVS results are highlighted in the red bouding box. The results demonstrate that our method consistently excels in reconstructing fine details and maintaining high color accuracy compared to other methods [26, 13, 29, 35, 30, 52, 3, 6].
Figure 13: Visualization of DVS and NVS of results. The DVS results are highlighted in the red bouding box. The results demonstrate that our method consistently excels in reconstructing fine details and maintaining high color accuracy compared to other methods.
@@
Appendix B Detailed Experiments
B.1 Synthetic Data Experiments
B.1.1 Deblurring View Synthesis Comparison
Refering to [41], we additionally provide deblurring view synthesis (DVS) results on our proposed EvaGaussians-Blender dataset, and show more qualitative results of novel view synthesis (NVS). For object-level scenes, Table. 5 and Figure. 12 present the quantitative and qualitative results of ours and the comparison baselines across six synthetic scene sequences. From the qualitative results, it is evident that our method excels in reconstructing fine details and maintaining high fidelity in both NVS and DVS. In terms of quantitative results, our method outperforms baseline methods in most scenes. For medium-scale scenes and large-scale scenes, the quantitative results are shown in Table. 6 and Table. 7, and the qualitative results of NVS and DVS are shown in Figure. 13, which demonstrate that our model achieves better performance in both tasks.
B.1.2 Per-scene Comparison for Novel View Synthesis
In this subsection, we present a detailed per-scene analysis of the novel view synthesis performance in medium and large scale scenes from the EvaGaussians-Blender dataset, to evaluate the effectiveness of our method across different challenging scenes.
Table. 8 shows the PSNR value of the NVS results, which demonstrates that our proposed method consistently outperforms other approaches across various scenes. The detailed metrics for SSIM and LPIPS in Table. 9 and Table. 10 further show that our model excels in maintaining structural integrity and perceptual quality in synthesized views. Specifically, in medium-scale scenes, our method exhibits robust performance, particularly in complex environments where maintaining detail and minimizing artifacts are challenging. This can be demonstrated in scenes such as cozyroom and factory, where our method achieves significant improvements in both PSNR and SSIM. For large-scale scenes, scenes like desert and city blocks highlight the model’s capability to generalize across different scales and provide high-quality novel view synthesis.
We also present more qualitative results of novel view synthesis in Figure. 14 and Figure. 15. The results highlight our model’s ability to reconstruct fine details and maintain high color accuracy beyond the comparison baselines.
B.2 Real-world Data Experiments
In this section, we present a comprehensive per-scene analysis of NVS results on the EvaGaussians-DAVIS dataset. The qualitative results are shown in Figure. 11, where the first column shows the blurry image used for training, and the following rows show the deblur results of different methods. The results demonstrate that our method consistently excels in reconstructing fine details compared to other methods.
We further report the per-scene quantitative results to validate our robustness to different scenes. As introudced in the main text, the adopted metrics include BRISQUE, NIQE, PIQE, MetaIQA, and RankIQA, which can effectively assess the quality of synthesized views in a no-reference manner. As shown in Table. 11, our model achieves the best BRISQUE scores across all scenes, highlighting its ability to produce visually appealing and less distorted images. For NIQE, as presented in Table. 13, our approach significantly outperforms the baselines, achieving the lowest average NIQE score. This demonstrates our method’s robustness in generating high-quality images with minimal perceptual artifacts. In terms of PIQE, Table. 14 shows that our model again leads in performance, achieving the lowest PIQE scores, which underscores the effectiveness of our model in preserving image details and reducing noise. Furthermore, our method excels in MetaIQA and RankIQA evaluations, as detailed in Tables. 15 and Table. 16, respectively. The highest MetaIQA scores and lowest RankIQA scores across most scenes affirm the overall better visual quality and fidelity of our synthesized views compared to baseline models. Overall, these results demonstrate the robustness of our method, particularly in handling complex scenes and maintaining high visual quality across diverse scenarios.
B.3 Ablation Study
In this section, we present an additional ablation study on the robustness of pose optimization in different blur level. We redesign three different levels of motion blur sequences in medium-scale scenes and compare the Average Trajectory Error (ATE) between the initial poses produced by COLMAP and the optimized poses. Figure. 7 illustrates the visualization of COLMAP poses and the optimized poses on the City Blocks scene. As the motion blur becomes more severe, the accuracy of the COLMAP poses is significantly impacted, while the optimized poses maintain a higher level of accuracy. In a horizontal comparison, the optimized poses better match the ground truth across various levels of blur, demonstrating the effectiveness of pose optimization. Table. 12 presents the quantitative results, further demonstrating the effectiveness and robustness of pose optimization in handling different levels of motion blur.
Appendix C Broader Impacts
Our proposed EvaGaussians leverages event cameras to assist novel view synthesis from low-quality, blurred images. It has the potential to bring about both positive and negative societal impacts.